OpenStudy (spring98):

Which expression is equivalent to 3^2 • 3^–5? A. 1/3^3 B. 1/3^7 C. 1/3^-3 D. 1/3^-7

jimthompson5910 (jim_thompson5910):

Use the rule \[\LARGE x^{\color{red}{a}}*x^{\color{blue}{b}} = x^{\color{red}{a}+\color{blue}{b}}\]

jimthompson5910 (jim_thompson5910):

For example \[\LARGE x^{\color{red}{2}}*x^{\color{blue}{3}} = x^{\color{red}{2}+\color{blue}{3}}=x^{5}\]

OpenStudy (spring98):

so is it C?

jimthompson5910 (jim_thompson5910):

what do you get when you use that rule (ignore the answer choices right now)

OpenStudy (spring98):

-3

jimthompson5910 (jim_thompson5910):

write out the whole thing, not just the exponent

OpenStudy (spring98):



jimthompson5910 (jim_thompson5910):

I meant to include the base

OpenStudy (spring98):

why?

jimthompson5910 (jim_thompson5910):

look back at the rule I posted

OpenStudy (spring98):

yes it's addition

OpenStudy (spring98):

i know

jimthompson5910 (jim_thompson5910):

so what does \(\LARGE 3^2*3^{-5}\) turn into when you use that rule?

OpenStudy (spring98):

3^-3

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

next comes the rule \[\LARGE x^{-a} = \frac{1}{x^a}\]

OpenStudy (spring98):

ohh so it's 1/3^-3

OpenStudy (spring98):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

do you see in the last rule how the `-a` exponent turned into `a` (without the negative) ?

OpenStudy (spring98):

yes

jimthompson5910 (jim_thompson5910):

so basically that rule is saying "if the exponent is negative, you take the reciprocal of the base to make the exponent positive"

jimthompson5910 (jim_thompson5910):

example \[\LARGE 2^{-7} = \frac{1}{2^7}\]

OpenStudy (spring98):

ok so that means it is not negative anymore so that the answer will be 1/3^3 right?

jimthompson5910 (jim_thompson5910):

yes, \[\LARGE 3^{-3} = \frac{1}{3^3}\]

OpenStudy (spring98):

thank you very much!!!

jimthompson5910 (jim_thompson5910):

no problem